The intensity of a certain siren is 0.060W/m^2 from a distance of 25 meters. What is the intensity of this siren at a distance of 50 meters? Answer in W/m^2
1
Expert's answer
2020-03-10T11:17:40-0400

The intensity of sound is inversely proportional to the square of distance between source and the observer.

"\frac{I_2}{I_1}=\frac{r_1^2}{r_2^2}=\frac{25^2}{50^2}=0.25"

"I_2=0.25I_1=0.25(0.06)=0.015\frac{W}{m^2}"